OpenStudy (anonymous):

Mr. Jimenez deposited money into an account in which interest is compounded quarterly at a rate of 2.6%. How much did he deposit if the total amount in his account after 4 years was $7160.06, and he made no other deposits or withdrawals?

OpenStudy (ybarrap):

In 4 years there are 4*4 = 16 quarters. Since each quarter, his principle increases by 1.026 % after 16 quarters, he has $$ \large{ P_0\times1.026^{16}=7,160.06 } $$ Where his initial deposit was \(P_0\). So, $$ \large{ P_o=\cfrac{7,160.06}{1.026^{16}} } $$ Does this make sense?

OpenStudy (anonymous):

A little confusing :P so what process am i suppose to go through to get the answer ?

OpenStudy (ybarrap):

We start with \(P_0\) dollars. At the end of the 1st quarter we have \(1.026\times P_0\) dollars. Are you with me?

OpenStudy (anonymous):

okay yes

OpenStudy (ybarrap):

In the 2nd quarter we have this and more $$ (1.026 \times P_0)\times 1.026 $$ Got that?

OpenStudy (ybarrap):

See where we are going?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

Ok, now. For ever quarter, we are multiplying by 1.026, because this represents the increase in the value of our previous amount. So after n quarters we have: $$ \large P_0\times1.026^n $$ Right?

OpenStudy (ybarrap):

For example, after 2 quarters, the value from his initial investment is $$ \large P_0\times1.026^2 $$

OpenStudy (ybarrap):

Because this equals $$ (1.026 \times P_0)\times 1.026 $$ As we discussed above.

OpenStudy (anonymous):

would i do this for 12 quarters?

OpenStudy (ybarrap):

In 4 years there are 16 quarters? Do you agree?

OpenStudy (anonymous):

right! so i would do that 16 times

OpenStudy (ybarrap):

You would do this 16 times, yes. Do it 16 times means $$ \large{ P_0\times 1.026^{16} } $$ Because, n, the number of quarters is 16 in 4 years. We are compounding every quarter.

OpenStudy (ybarrap):

We know that value of our investment after 16 quarters, right? So this number we just developed must equal 7160.06

OpenStudy (ybarrap):

The only thing left for us to do is to find our initial investment \(P_0\). Do you have this?

OpenStudy (ybarrap):

Are we in sync?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

k, so let's find \(P_0\). Can you help me?

OpenStudy (anonymous):

would it be 4748.53

OpenStudy (ybarrap):

I don't know, what did you do? Tell me.

OpenStudy (ybarrap):

Did you use our work here?

OpenStudy (ybarrap):

Whatever this turns out to be is the answer... $$ \large{ P_o=\cfrac{7,160.06}{1.026^{16}} } $$

OpenStudy (anonymous):

yes thats what i did

OpenStudy (ybarrap):

lemme double check for you

OpenStudy (ybarrap):

Yes maam! You are correcto!

OpenStudy (anonymous):

okay

OpenStudy (ybarrap):

great!